# Fast relaxation of a viscous vortex in an external flow

### v1

@@ 1 Introduction
Figure 1: Numerical simulation of a vortex in an external field with Gaussian initial data. The vorticity distribution (left) and the deviation from the Lamb-Oseen vortex (right) are represented at nine different times, using standard color codes for the vorticity levels. The final state at $t=7$ is close to the approximate solution defined in (1.10). This simulation is made with the free software Basilisk, and the external field is chosen as in Appendix A.1.
The purpose of this paper is twofold. First, we show that the techniques introduced in [4] to study the solution of the two-dimensional Navier-Stokes equations with a finite collection of point vortices as initial data can be adapted to the emblematic case of a single vortex in an external flow, which is at the same time simpler and more general. In particular, if the initial vorticity is a Dirac mass, we construct perturbatively an accurate approximation of the solution, and we verify that the exact solution remains close to it over a long time interval if the viscosity is small enough. Next, we consider ill-prepared data for which the initial vorticity is a sharply concentrated Gaussian function, and we prove that the resulting solution rapidly relaxes towards the approximate solution computed in the well-prepared case. That part of the analysis relies on enhanced dissipation estimates for the linearized Navier-Stokes equations at the Lamb-Oseen vortex, which are due to Li, Wei, and Zhang [13]. Such estimates were already applied in [5] to quantify the stability properties of a Gaussian vortex in the regime of high Reynolds numbers, but to our knowledge they were never used to study the deformation of vortices under external strain.
@@
where the parameter $\nu>0$ is the kinematic viscosity of the the fluid. The velocity field $u=(u_{1},u_{2})$ associated with $\omega$ is given by the Biot-Savart formula
$u(x,t)\,=\,\frac{1}{2\pi}\int_{\mathbb{R}^{2}}\frac{(x-y)^{\perp}}{|x-y|^{2}}\,\omega(y,t)\,\mathrm{d}y\,,\qquad\qquad\forall\,(x,t)\in\mathbb{R}^{2}\times(0,T)\,,$ (1.3)
@@
Equation (1.2) appears in at least two physical contexts. The first one is the evolution of a finite number of isolated vortices under the dynamics of the Navier-Stokes equations in $\mathbb{R}^{2}$. A natural strategy in this situation is to concentrate on the motion of one particular vortex, in which case the corresponding vorticity distribution $\omega$ satisfies the evolution equation (1.2) with $u=\mathrm{BS}[\omega]$ and $f$ being the velocity field associated with all the other vortices in the flow, see for example [19, 10, 4]. Note that the vorticity $\omega_{f}:=\partial_{x_{1}}f_{2}-\partial_{x_{2}}f_{1}$ is advected by the velocity field $u$, so that the external flow $f$ is not independent of the solution $\omega$ of (1.2). Alternatively, following [24, 25], we can consider the evolution of a single vortex in a background potential flow $f$, which is typically due to an inflow condition at infinity. In that case $\omega_{f}=0$, so that the dynamics of the vortex does not influence the external flow.
@@
Identifying the core size $\ell$ with the diffusion length $\sqrt{\nu t}$, we see that the first term in the right-hand side of (1.10) is the exactly Lamb-Oseen vortex (1.5), which is a radially symmetric function of the rescaled variable $(x-z)/\ell$; in contrast, the correction term involving $w_{2}$ depends explicitly on the polar angle $\theta$. In view of (1.1) the strain rates (1.11) are bounded by $T_{0}^{-1}$, so that
@@
The most natural way of locating the position of a concentrated vortex with nonzero circulation $\Gamma$ is to use the center of vorticity $\bar{z}(t)$, which satisfies
@@
Under the assumptions of Theorem 1.6, we show in Section 3.3.4 that $|\bar{z}(t)-z(t)|\leq Cd\varepsilon^{3}\bigl{(}\varepsilon+\delta\bigr{)}$ for some constant $C>0$. This means that the motion of the center of vorticity is accurately described by the ODE (1.14), and that estimate (1.13) still holds if $z(t)$ is replaced by $\bar{z}(t)$. In contrast, using the approximate vortex position $\hat{z}(t)$ leads to a less precise control on the solution, as in (1.7).
@@
Since $\beta=c_{2}\delta^{-1/3}$ we have $(t_{0}/t)^{\beta}\leq\delta$ when $t\geq t_{0}(1+\tau_{\delta})$, where $\tau_{\delta}=c_{3}\delta^{1/3}\log(1/\delta)$ for some $c_{3}>0$. The right-hand side of (1.17) is therefore of size $\varepsilon(t)^{2}\delta^{1/6}\bigl{(}\log(1/\delta)\bigr{)}^{1/2}$ as soon as $t\geq t_{0}(1+\tau_{\delta})$. In other words, the solution of (1.2) rapidly relaxes towards the approximate solution (1.10), which takes into account the effect of the external strain, and remains close to it up to the final time $T$. This description agrees with the numerical observation in Figure 1. That the relaxation rate $\beta$ depends on the inverse Reynolds number $\delta=\nu/\Gamma$ is a consequence of the enhanced dissipation effect in the vortex core, see [5] and Section 4.
@@
The proof of our results relies on the construction of an approximate solution of the initial value problem in self-similar variables, which is carried out in Section 2. This part of the argument closely follows the previous works [4, 3] where particular situations were considered. The proof of Theorem 1.6 is carried out in Section 3, first under the simplifying assumption that $T/T_{0}\ll 1$, and then for any $T>0$. In both cases the desired control on the solution is obtained by an energy estimate in some weighted $L^{2}$ space, but the construction of the weight function is much more complicated if $T$ is not small compared to $T_{0}$. Improving upon the results of [4], we construct a Gaussian-like weight that provides an optimal control on the solution as far as the decay at infinity is concerned, and implies in particular the $L^{1}$ estimate (1.13). In Section 4, we show how these arguments can be combined with the enhanced dissipation estimates obtained by Li, Wei, and Zhang [13] to yield a proof of Theorem 1.8. Finally, a few auxiliary results are collected in the Appendix. In particular, we clarify the link between our approximate solution (1.10) and the Burgers vortex in an asymmetric strain, and we investigate the motion of the center of vorticity under the assumptions of Theorem 1.6.
@@ 2 Self-similar variables and approximate solution
The new space variable $\xi$ measures the distance to the vortex center $z(t)$ in units of the diffusion length $\sqrt{\nu t}$. As already explained, the small parameter $\delta$ is the inverse Reynolds number, and the time-dependent aspect ratio $\varepsilon$ compares the size $\sqrt{\nu t}$ of the vortex core to the effective size $d$ of the vortex.
@@
The position $z(t)$ of the vortex center is unknown at this stage, but will be chosen so as to minimize the quantity $f(z(t)+\sqrt{\nu t}\,\xi,t)-z^{\prime}(t)$ in an appropriate sense. The leading order approximation corresponds to the natural choice $z^{\prime}(t)=f(z(t),t)$, but higher order corrections, leading to the modified ODE (1.14), will be needed to achieve the desired precision. We also observe that the rescaled velocity $U(\xi,t)$ is divergence-free and satisfies $\partial_{1}U_{2}-\partial_{2}U_{1}=\Omega$, which means that $U$ is obtained from $\Omega$ by the Biot-Savart formula (1.3), namely $U=\mathrm{BS}[\Omega]$.
It is important to observe that (2.3) is not a regular evolution equation at time $t=0$, due to the singular time derivative $t\partial_{t}\Omega$ in the left-hand side. Nevertheless, if we adapt to the present case the results of [8, 10], which hold for $f(z,t)=0$ and $z(t)=0$, it is not difficult to show that (2.3) has a unique (mild) solution $\Omega\in C^{0}((0,T],L^{1}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2}))$ that satisfies $\|\Omega(\cdot,t)-\Omega_{0}\|_{L^{1}}\to 0$ as $t\to 0$. This is precisely the solution we study in Theorem 1.6. Note that the Gaussian profile (1.5) is, up to normalization, the only possible choice for the initial vorticity at time $t=0$. The situation considered in Theorem 1.8 is much different: the Cauchy problem for equation (2.3) is well-posed at any positive positive time $t_{0}>0$, and we could therefore choose arbitrary initial data at $t=t_{0}$. However, for reasons that are explained in Remark 1.10 above, our choice is to take the same initial vorticity $\Omega_{0}$ as in Theorem 1.6.
@@ 2.1 Expansion of the external velocity
where $\xi$ is the polar angle of the variable $\xi\in\mathbb{R}^{2}$, and $a(t),b(t),c(t)$ denote the following derivatives of $f$ evaluated at $(z(t),t)$:
@@
where $\xi$ is the polar angle of the variable $\xi\in\mathbb{R}^{2}$, and $A=\frac{3}{8}\partial_{1}^{2}f_{1}-\frac{1}{8}\partial_{2}^{2}f_{1}$, $B=\frac{1}{8}\partial_{1}^{2}f_{2}-\frac{3}{8}\partial_{2}^{2}f_{2}$. All derivatives of $f$ are evaluated at $(z(t),t)$.
@@ 2.2 Functional framework
where $\Omega_{0},U_{0}$ are given by (1.6). If we consider the operators $\mathcal{L},\Lambda$ as acting on the function space (2.11), with maximal domain, we have the following result:
@@ 2.3.1 Second order vorticity profile
such that $\varphi_{2}(r)=\mathcal{O}(r^{2})$ as $r\to 0$ and $\phi(r)=\mathcal{O}(r^{-2})$ as $r\to+\infty$. In particular $w_{2}(r)>0$ for all $r>0$, $w_{2}(r)=\mathcal{O}(r^{2})$ as $r\to 0$, and $w_{2}(r)\sim(r^{4}/8)e^{-r^{2}/4}$ as $r\to+\infty$.
@@ 2.3.3 Fourth order vorticity profile
$\mathcal{R}_{4}\,=\,\delta(2-\mathcal{L})\Omega_{4}+\delta\bigl{(}\bar{U}_{2}+E_{2}\bigr{)}\cdot\nabla\tilde{\Omega}_{2}+\delta\tilde{U}_{2}\cdot\nabla\Omega_{2}+T_{0}\partial_{t}\tilde{\Omega}_{2}\,=\,\mathcal{O}(\delta)\,.$ (2.31)
@@ 3 The solution starting from a point vortex
In this section we complete the proof of Theorem 1.6. We assume throughout that the kinematic viscosity $\nu>0$ is small compared to the circulation parameter $\Gamma>0$, which is fixed once and for all. Given any $z_{0}\in\mathbb{R}^{2}$, we consider the unique solution $\omega(x,t)$ of the vorticity equation (1.2), (1.3) satisfying the conditions (1.4), which imply that the initial vorticity is a Dirac mass of strength $\Gamma$ located at point $z_{0}$. Since (1.2) is a viscous conservation law, we know in particular that $\int_{\mathbb{R}^{2}}\omega(x,t)\,\mathrm{d}x=\Gamma$ for all positive times. To desingularize the solution in the regime where $\nu t$ is small, we make the change of variables (2.1), where $z(t)$ denotes the unique solution of the modified ODE (1.14) such that $z(0)=z_{0}$. The rescaled vorticity $\Omega(\xi,t)$ and the associated velocity $U(\xi,t)$ then satisfy the evolution equation (2.6) with initial data (1.6).
@@
for all $t\in(0,T)$ and all $\xi\in\mathbb{R}^{2}$, where $\Omega_{\mathrm{app}}$ is the approximate solution (2.5) and $U_{\mathrm{app}}$ is the associated velocity field. By construction, the correction terms $w(\xi,t),v(\xi,t)$ in (3.1) vanish at initial time $t=0$, and our goal is to show that they remain small in an appropriate topology for all $t\in[0,T]$. In view of (2.6), the vorticity $w(\cdot,t)$ satisfies the evolution equation
@@
for $t\in(0,T)$, where $\delta=\nu/\Gamma$ and $\mathcal{R}_{\mathrm{app}}$ is the remainder term (2.21). The velocity $v(\cdot,t)$ is obtained from $w(\cdot,t)$ by the usual Biot-Savart formula. Since $\int\Omega(\xi,t)\,\mathrm{d}\xi=1$, it follows from Remark 2.10 that
@@
The main difficulty in our analysis is the necessity of controlling the solution of (3.2) uniformly in the small parameter $\delta>0$. Since we consider zero initial data, the evolution is entirely driven by the source term $\delta^{-2}\mathcal{R}_{\mathrm{app}}$ in the right-hand side, which is of size $\mathcal{O}(\varepsilon^{2}+\delta^{-2}\varepsilon^{5})$ according to Proposition 2.9. Note that the fifth power of $\varepsilon$ in estimate (2.32) is due to our choice of constructing the approximate solution (2.5) as a fourth order expansion in $\varepsilon$. Now, it follows from the definitions (2.2) that $\delta^{-1}\varepsilon^{2}=t/T_{0}$, which means that $\delta^{-2}\varepsilon^{5}=\mathcal{O}(\varepsilon)$ as long as $t$ is comparable with $T_{0}$.
To prove that the solution of (3.2) remains of size $\mathcal{O}(\varepsilon)$ over the whole time interval $[0,T]$, we have to show that the contributions of the source term do not get excessively amplified by the linear terms in (3.2), which are multiplied by the large factor $\delta^{-1}$. This can be done using an appropriate energy estimate in a weighted $L^{2}$ space, where the weight function is carefully chosen so as to minimize the contributions of the dangerous linear terms in (3.2). Note that the nonlinear term $v\cdot\nabla w$ is not multipled by $\delta^{-1}$, because we chose to include a factor of $\delta$ in the definition (3.1) of the corrections terms $w,v$.

### v2

@@ 1 Introduction
Figure 1: Numerical simulation of a vortex in an external field with Gaussian initial data. The vorticity distribution (left) and the deviation from the Lamb-Oseen vortex (right) are represented at nine different times, using standard color codes for the vorticity levels. The final state at $t=7$ is close to the approximate solution defined in (1.10). This simulation is made with the free software Basilisk, and the external field is chosen as in Section A.1.
The purpose of this paper is twofold. First, we show that the techniques introduced in [4] to study the solution of the two-dimensional Navier-Stokes equations with a finite collection of point vortices as initial data can be adapted to the emblematic case of a single vortex in an external flow, which is at the same time simpler and more general. In particular, if the initial vorticity is a Dirac mass, we construct perturbatively an accurate approximation of the solution, and we verify that the exact solution remains close to it over a long time interval if the viscosity is small enough. Next, we consider ill-prepared data for which the initial vorticity is a sharply concentrated Gaussian function, and we prove that the resulting solution rapidly relaxes towards the approximate solution computed in the well-prepared case. That part of the analysis relies on enhanced dissipation estimates for the linearized Navier-Stokes equations at the Lamb-Oseen vortex, which are due to Li, Wei, and Zhang [13]. Such estimates were already applied in [5] to prove axisymmetrization near a Gaussian vortex in the regime of high Reynolds numbers, but to our knowledge they were never used to study the relaxation of a circular vortex towards a non-symmetric metastable state in an external strain.
@@
where the parameter $\nu>0$ is the kinematic viscosity of the fluid. The velocity field $u=(u_{1},u_{2})$ associated with $\omega$ is given by the Biot-Savart formula
$u(x,t)\,=\,\frac{1}{2\pi}\int_{\mathbb{R}^{2}}\frac{(x-y)^{\perp}}{|x-y|^{2}}\,\omega(y,t)\,\mathrm{d}y\,,\qquad\forall\,(x,t)\in\mathbb{R}^{2}\times(0,T)\,,$ (1.3)
@@
Equation (1.2) appears in at least two physical contexts. The first one is the evolution of a finite number of isolated vortices under the dynamics of the Navier-Stokes equations in $\mathbb{R}^{2}$. A natural strategy in this situation is to concentrate on the motion of one particular vortex, whose vorticity distribution $\omega$ satisfies the evolution equation (1.2) where $u=\mathrm{BS}[\omega]$ and $f$ is the velocity field associated with the other vortices in the flow, see for example [19, 10, 4]. In that case the vorticity $\omega_{f}:=\partial_{x_{1}}f_{2}-\partial_{x_{2}}f_{1}$ is advected by the velocity field $u$, so that the external flow $f$ is not independent of the solution $\omega$ of (1.2). Alternatively, following [24, 25], we can consider the evolution of a single vortex in a background potential flow $f$, which is typically due to an inflow condition at infinity. Here $\omega_{f}=0$, so that the dynamics of the vortex does not influence the external flow.
@@
Identifying the core size $\ell$ with the diffusion length $\sqrt{\nu t}$, we see that the first term in the right-hand side of (1.10) is exactly the Lamb-Oseen vortex (1.5), which is a radially symmetric function of the rescaled variable $(x-z)/\ell$; in contrast, the correction term involving $w_{2}$ depends explicitly on the polar angle $\theta$. In view of (1.1) the strain rates (1.11) are bounded by $T_{0}^{-1}$, so that
@@
The most natural way of locating the position of a concentrated vortex that evolves according to (1.2) is to use the center of vorticity $\bar{z}(t)$, which satisfies
@@
Under the assumptions of Theorem 1.6, we show in Section 3.3.4 that $|\bar{z}(t)-z(t)|\leq Cd\varepsilon^{3}\bigl{(}\varepsilon+\delta\bigr{)}$ for some constant $C>0$. This means that the motion of the center of vorticity is accurately described by the ODE (1.14), and that estimate (1.13) still holds if $z(t)$ is replaced by $\bar{z}(t)$. In contrast, using the naive vortex position $\hat{z}(t)$ deteriorates the precision of our approximate solution, as can be seen from the right-hand side of (1.7) which is $\mathcal{O}(\varepsilon)$ instead of $\mathcal{O}(\varepsilon^{2})$.
@@
Since $\beta=c_{2}\delta^{-1/3}$ we have $(t_{0}/t)^{\beta}\leq\delta$ when $t\geq t_{0}(1+\tau_{\delta})$, where $\tau_{\delta}=c_{3}\delta^{1/3}\log(1/\delta)$ for some $c_{3}>0$. The right-hand side of (1.17) is therefore of size $\varepsilon(t)^{2}\delta^{1/6}\bigl{(}\log(1/\delta)\bigr{)}^{1/2}$ as soon as $t\geq t_{0}(1+\tau_{\delta})$. In other words, the solution of (1.2) rapidly relaxes towards the approximate solution (1.10), which takes into account the effect of the external strain, and remains close to it up to the final time $T$. This description agrees with the numerical observations in Figure 1. That the relaxation rate $\beta$ depends on the inverse Reynolds number $\delta=\nu/\Gamma$ is a consequence of the enhanced dissipation effect in the vortex core, see [5] and Section 4.
@@
The proof of our results relies on the construction of an approximate solution of the initial value problem in self-similar variables, which is performed in Section 2. This part of the argument closely follows the previous works [4, 3] where particular situations were considered. The proof of Theorem 1.6 is carried out in Section 3, first under the simplifying assumption that $T/T_{0}\ll 1$, and then for any $T>0$. In both cases the desired control on the solution is obtained by an energy estimate in some weighted $L^{2}$ space, but the construction of the weight function is much more complicated if $T$ is not small compared to $T_{0}$. Improving upon the results of [4], we construct a Gaussian-like weight that provides an optimal control on the solution as far as the decay at infinity is concerned, and implies in particular the $L^{1}$ estimate (1.13). In Section 4, we show how these arguments can be combined with the enhanced dissipation estimates obtained by Li, Wei, and Zhang [13] to yield a proof of Theorem 1.8. Finally, a few auxiliary results are collected in the Appendix. In particular, we clarify the link between our approximate solution (1.10) and the Burgers vortex in an asymmetric strain, and we investigate the motion of the center of vorticity under the assumptions of Theorem 1.6.
@@ 2 Self-similar variables and approximate solution
The new space variable $\xi$ describes the position with respect to the vortex center $z(t)$ measured in units of the diffusion length $\sqrt{\nu t}$. As already explained, the small parameter $\delta$ is the inverse Reynolds number, and the time-dependent aspect ratio $\varepsilon$ compares the size $\sqrt{\nu t}$ of the vortex core to the effective size $d$ of the vortex.
@@
The position $z(t)$ of the vortex center is unknown at this stage, but will be chosen so as to minimize the quantity $f(z(t)+\sqrt{\nu t}\,\xi,t)-z^{\prime}(t)$ in an appropriate sense. The natural choice $z^{\prime}(t)=f(z(t),t)$ gives the leading order approximation, but higher order corrections will be needed to achieve the desired precision. We also observe that the rescaled velocity $U(\xi,t)$ is divergence-free and satisfies $\partial_{1}U_{2}-\partial_{2}U_{1}=\Omega$, which means that $U$ is obtained from $\Omega$ by the Biot-Savart formula (1.3), namely $U=\mathrm{BS}[\Omega]$.
It is important to realize that (2.3) is not a regular evolution equation at time $t=0$, due to the singular time derivative $t\partial_{t}\Omega$ in the left-hand side. Nevertheless, if we adapt to the present case the results of [8, 10], which hold for $f(z,t)=0$ and $z(t)=0$, it is not difficult to show that (2.3) has a unique (mild) solution $\Omega\in C^{0}((0,T],L^{1}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2}))$ that satisfies $\|\Omega(\cdot,t)-\Omega_{0}\|_{L^{1}}\to 0$ as $t\to 0$. This is precisely the solution we study in Theorem 1.6. Note that the Gaussian profile (1.5) is, up to normalization, the only possibility for the initial vorticity at time $t=0$. The situation considered in Theorem 1.8 is much different: the Cauchy problem for equation (2.3) is well-posed at any positive positive time $t_{0}>0$, and we could therefore choose arbitrary initial data at $t=t_{0}$. However, for reasons that are explained in Remark 1.10 above, our choice is to take the same initial vorticity $\Omega_{0}$ as in Theorem 1.6.
@@ 2.1 Expansion of the external velocity
where $\theta$ is the polar angle of the variable $\xi\in\mathbb{R}^{2}$, and $a(t),b(t),c(t)$ denote the following derivatives of $f$ evaluated at $(z(t),t)$:
@@
where $\theta$ is the polar angle of the variable $\xi\in\mathbb{R}^{2}$, and $A=\frac{3}{8}\partial_{1}^{2}f_{1}-\frac{1}{8}\partial_{2}^{2}f_{1}$, $B=\frac{1}{8}\partial_{1}^{2}f_{2}-\frac{3}{8}\partial_{2}^{2}f_{2}$. All derivatives of $f$ are evaluated at $(z(t),t)$.
@@ 2.2 Functional framework
where $\Omega_{0},U_{0}$ are given by (1.6). If we consider the operators $\mathcal{L},\Lambda$ as acting on the function space (2.11) with maximal domain, we have the following result:
@@ 2.3.1 Second order vorticity profile
such that $\varphi_{2}(r)=\mathcal{O}(r^{2})$ as $r\to 0$ and $\phi(r)=\mathcal{O}(r^{-2})$ as $r\to+\infty$. In particular $w_{2}(r)>0$ for all $r>0$, $w_{2}(r)=\mathcal{O}(r^{2})$ as $r\to 0$, and $w_{2}(r)\sim(r^{4}/8)e^{-r^{2}/4}$ as $r\to+\infty$. The graph of the function $w_{2}$ is represented in Figure 2.
@@ 2.3.3 Fourth order vorticity profile
$\mathcal{R}_{4}\,=\,\delta(2-\mathcal{L})\Omega_{4}+\delta\bigl{(}\bar{U}_{2}+E_{2}\bigr{)}\cdot\nabla\tilde{\Omega}_{2}+\delta\tilde{U}_{2}\cdot\nabla\Omega_{2}+\delta T_{0}\partial_{t}\tilde{\Omega}_{2}\,=\,\mathcal{O}(\delta)\,.$ (2.31)
@@ 3 The solution starting from a point vortex
In this section we complete the proof of Theorem 1.6. We assume throughout that the kinematic viscosity $\nu>0$ is small compared to the circulation parameter $\Gamma>0$, which is fixed once and for all. Given any $z_{0}\in\mathbb{R}^{2}$, we consider the unique solution $\omega(x,t)$ of the vorticity equation (1.2), (1.3) satisfying the conditions (1.4), which imply that the initial vorticity is a Dirac mass of strength $\Gamma$ located at point $z_{0}$. Since (1.2) is a viscous conservation law, we know in particular that $\int_{\mathbb{R}^{2}}\omega(x,t)\,\mathrm{d}x=\Gamma$ for all positive times. To desingularize the solution in the regime where $\nu t$ is small, we make the change of variables (2.1), where $z(t)$ denotes the unique solution of the ODE (1.14) such that $z(0)=z_{0}$. The rescaled vorticity $\Omega(\xi,t)$ and the associated velocity $U(\xi,t)$ then satisfy the evolution equation (2.6) with initial data (1.6).
@@
for all $t\in(0,T)$ and all $\xi\in\mathbb{R}^{2}$, where $\Omega_{\mathrm{app}}$ is the approximate solution (2.5), $U_{\mathrm{app}}$ is the associated velocity field, and $\delta=\nu/\Gamma$. By construction, the correction terms $w(\xi,t),v(\xi,t)$ in (3.1) vanish at initial time $t=0$, and our goal is to show that they remain small in an appropriate topology for all $t\in[0,T]$. In view of (2.6), the vorticity $w(\cdot,t)$ satisfies the evolution equation
@@
for $t\in(0,T)$, where $\mathcal{R}_{\mathrm{app}}$ is the remainder term (2.21). The velocity $v(\cdot,t)$ is obtained from the vorticity $w(\cdot,t)$ by the usual Biot-Savart formula (1.3). Since $\int\Omega(\xi,t)\,\mathrm{d}\xi=1$, it follows from Remark 2.10 that
@@
The main difficulty in our analysis is the necessity of controlling the solution of (3.2) uniformly in the small parameter $\delta>0$. Since we consider zero initial data, the evolution is entirely driven by the source term $\delta^{-2}\mathcal{R}_{\mathrm{app}}$ in the right-hand side, which is of size $\mathcal{O}(\varepsilon^{2}+\delta^{-2}\varepsilon^{5})$ according to Proposition 2.9. Here and in (2.32), the fifth power of $\varepsilon$ is due to our choice of constructing the approximate solution (2.5) as a fourth order expansion in $\varepsilon$. This is sufficient to counterbalance the large factor $\delta^{-2}$ because $\delta^{-1}\varepsilon^{2}=t/T_{0}$ by (2.2), which means that $\delta^{-2}\varepsilon^{5}=\mathcal{O}(\varepsilon)$ as long as $t$ remains comparable with $T_{0}$.
To prove that the solution of (3.2) remains of size $\mathcal{O}(\varepsilon)$ over the whole time interval $[0,T]$, we have to show that the linear terms in (3.2), which are multiplied by the large factor $\delta^{-1}$, do not create instabilities that could result in a rapid amplification of the solution, on a timescale proportional to $\delta$. This can be done using an appropriate energy estimate in a weighted $L^{2}$ space, where the weight function is carefully chosen so as to minimize the contributions of the dangerous linear terms. Note that the nonlinearity $v\cdot\nabla w$ in (3.2) is not multipled by $\delta^{-1}$, because we chose to include a factor of $\delta$ in the definition (3.1) of the corrections terms $w,v$.